# Bundle Neural Networks for message diffusion on graphs

### v1

@@ Abstract
The dominant paradigm for learning on graph-structured data is message passing. Despite being a strong inductive bias, the local message passing mechanism suffers from pathological issues such as over-smoothing, over-squashing, and limited node-level expressivity. To address these limitations we propose Bundle Neural Networks (BuNN), a new type of GNN that operates via message diffusion over flat vector bundles – structures analogous to connections on Riemannian manifolds that augment the graph by assigning to each node a vector space and an orthogonal map. A BuNN layer evolves the features according to a diffusion-type partial differential equation. When discretized, BuNNs are a special case of Sheaf Neural Networks (SNNs), a recently proposed MPNN capable of mitigating over-smoothing. The continuous nature of message diffusion enables BuNNs to operate on larger scales of the graph and, therefore, to mitigate over-squashing. Finally, we prove that BuNN can approximate any feature transformation over nodes on any (potentially infinite) family of graphs given injective positional encodings, resulting in universal node-level expressivity. We support our theory via synthetic experiments, and showcase the strong empirical performance of BuNNs over a range of real-world tasks, achieving state-of-the-art results on several standard benchmarks in transductive and inductive settings.
@@ 1 Introduction
Graph Neural Networks (GNNs) [1, 2] are widely adopted machine learning models designed to operate over graph structures. A majority of GNN models function locally, respecting the topology by having nodes exchange messages directly with neighboring nodes. Such models are referred to as Message Passing Neural Networks (MPNNs) [3]. While message passing has repeatedly proven to be a useful inductive bias in many tasks, the local mechanism was shown to suffer from certain pathological issues, namely over-smoothing [4, 5, 6], over-squashing [7, 8, 9], and limited expressivity [10, 11].
To address these issues, we propose Bundle Neural Networks (BuNNs), a new type of GNN that operates over flat vector bundles – a structure that augments the graph by assigning to each node a vector space and an orthogonal map. BuNNs do not perform ‘explicit’ message passing through multiple steps of information exchange between neighboring nodes, but instead operate via what we call message diffusion. Each layer involves a node update step, and a diffusion step evolving the features according to a vector diffusion PDE as in Singer and Wu [12]. The resulting architecture enjoys the desirable properties of Sheaf Neural Networks [13], in that they can avoid over-smoothing, but are more scalable and achieve better performance on a range of benchmark datasets. We also show how the continuous nature of message diffusion together with the bundle structure allows to mitigate issues such as over-squashing. Additionally, we prove that under the right conditions, BuNNs are compact uniform approximators, a type of universality result for node-level function approximation.
@@
  • We study the diffusion process and show that it can mitigate over-smoothing and over-squashing (Section 4). We support these claims with synthetic experiments (Section 6.1).
  • We prove that, under certain conditions, BuNNs are compact uniform universal approximators, and this is, to our knowledge, the first of such results (Section 5).
  • We show that BuNNs perform well on a range of tasks, for instance achieving a new state-of-the-art result on the Peptides-func dataset (Section 6.2).
@@ 2 Background
Let $\mathsf{G}=(\mathsf{V},\mathsf{E})$ be an undirected graph on $n=|\mathsf{V}|$ nodes with edges $\mathsf{E}$. We represent the edges via an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ where the entry $\mathbf{A}_{uv}$ for $u,v\in\mathsf{V}$ is $1$ if the edge $(u,v)\in\mathsf{E}$ and 00 otherwise. Let $\mathbf{D}$ be the diagonal degree matrix with entry $\mathbf{D}_{vv}=d_{v}$ equal to the degree of $v$. The graph Laplacian is defined as $\mathbf{L}:=\mathbf{D}-\mathbf{A}$ and the random walk normalized graph Laplacian is defined as $\mathbf{\mathcal{L}}:=\mathbf{I}-\mathbf{D^{-1}A}$. We assume that at each node $v\in\mathsf{V}$ we are given a $c$-dimensional signal $\mathbf{x}_{v}$ and group such signals into a matrix $\mathbf{X}\in\mathbb{R}^{n\times c}$. A feature transformation on a graph $\mathsf{G}$ is a permutation-equivariant map $f_{\mathsf{G}}:\mathbb{R}^{n\times c}\rightarrow\mathbb{R}^{n\times c^{\prime}}$ that transforms the node signals.
@@
GNNs and feature transformations. A $\operatorname{GNN}_{\boldsymbol{\Theta}}$ is a (continuous) map parameterized by $\boldsymbol{\Theta}$ that takes as input a graph alongside node signals $\mathsf{G}=(\mathsf{V},\mathsf{E},\mathbf{X})$ and outputs a transformed signal $(\mathsf{V},\mathsf{E},\mathbf{X^{\prime}})$. A GNN on a graph $\mathsf{G}$ is therefore a feature transformation $\operatorname{GNN}_{\Theta}:\mathbb{R}^{n\times c}\rightarrow\mathbb{R}^{n\times c^{\prime}}$. Given a collection of graphs $\mathcal{G}$, with feature space $\mathbb{R}$, a feature transformation $F$ of $\mathcal{G}$ is an assignment of every graph $\mathsf{G}\in\mathcal{G}$ to a feature transformation $F_{G}:\mathbb{R}^{n_{G}\times c}\rightarrow\mathbb{R}^{n_{G}\times c^{\prime}}$. The set of (continuous) feature transformations over a collection of graphs in $\mathcal{G}$ is denoted $\mathcal{C}\left(\mathcal{G},\mathbb{R}^{c},\mathbb{R}^{c^{\prime}}\right)$.
@@
Cellular sheaves. A cellular sheaf [14] $\left(\mathcal{F},\mathsf{G}\right)$ over an undirected graph $\mathsf{G}=\left(\mathsf{V},\mathsf{E}\right)$ augments $\mathsf{G}$ by attaching to each node $v$ and edge $e$ a vector space space called stalks and denoted by $\mathcal{F}(v)$ and $\mathcal{F}(e)$, usually the stalks are copies of $\mathbb{R}^{d}$ for some $d$. Additionally, every incident node-edge pair $v\trianglelefteq e$ gets assigned a linear map between stalks called restriction maps and denoted $\mathcal{F}_{v\trianglelefteq e}:\mathcal{F}(v)\rightarrow\mathcal{F}(e)$. Given two nodes $v$ and $u$ connected by an edge $(v,\ u)$, we can transport a vector $\mathbf{x}_{v}\in\mathcal{F}(v)$ from $v$ to $u$ by first mapping it to the stalk at $e=(v,u)$ using $\mathcal{F}_{v\trianglelefteq e}$, and mapping it to $\mathcal{F}(u)$ using the transpose $\mathcal{F}_{u\trianglelefteq e}^{T}$. As a generalization of the graph adjacency matrix, the sheaf adjacency matrix $\mathbf{A}_{\mathcal{F}}\in\mathbb{R}^{nd\times nd}$ is defined as a block matrix in which each $d\times d$ block $(\mathbf{A}_{\mathcal{F}})_{uv}$ is $\mathcal{F}_{u\trianglelefteq e}^{T}\mathcal{F}_{v\trianglelefteq e}$ if there is an edge between $u$ and $v$ and $\mathbf{0}_{d\times d}$ if there is not. We similarly define the block diagonal degree matrix $\mathbf{D}_{\mathcal{F}}\in\mathbb{R}^{nd\times nd}$ as $\left(\mathbf{D}_{\mathcal{F}}\right)_{vv}:=d_{v}\mathbf{I}_{d\times d}$. Finally, we define the sheaf Laplacian as $\mathbf{L}_{\mathcal{F}}=\mathbf{D_{\mathcal{F}}}-\mathbf{A}_{\mathcal{F}}$. These matrices act as bundle generalizations of their well-known standard graph counterparts and we recover such matrices when $\mathcal{F}(v)\cong\mathbb{R}$ and $\mathcal{F}_{v\trianglelefteq e}{v}=1$ for all $v\in\mathsf{V}$ and $e\in\mathsf{E}$.
@@
Vector bundles. When restriction maps are orthogonal, we call the sheaf a vector bundle, a structure analogous to connections on Riemannian manifolds. For this reason, the sheaf Laplacian also takes the name connection Laplacian [12]. In this case the product $\mathcal{F}_{u\trianglelefteq e}^{T}\mathcal{F}_{v\trianglelefteq e}$ is also orthogonal and is denoted $\mathbf{O}_{uv}$ referring to the transformation a vector undergoes when moved across a manifold via parallel transport. In this case we denote the node-stalk at $v$ by $\mathcal{B}(v)$, the bundle-adjacency by $\mathbf{A}_{\mathcal{B}}$ and the bundle Laplacian $\mathbf{L}_{\mathcal{B}}$, and its random-walk normalized version $\mathbf{\mathcal{L}}_{\mathcal{B}}:=\mathbf{I}_{dn\times dn}-\mathbf{D}_{\mathcal{B}}^{-1}\mathbf{A}_{\mathcal{B}}$.
@@ 3 Bundle Neural Networks
In this Section, we derive the Bundle Neural Network model from heat diffusion equations over flat vector bundles. We then discuss its relationship to GCN [15] and NSD [13]. We provide algorithmic implementations and additional practical details in the Appendix (Section E).
@@
The gradient of the Dirichlet energy $\nabla_{\mathbf{X}}\mathcal{E}_{\mathcal{B}}(\mathbf{X})$ is the random-walk Laplacian $\mathbf{\mathcal{L}}_{\mathcal{B}}$. We can write down a heat diffusion equation (or simply heat equation) over a vector bundle as a gradient flow, whose evolution equation with initial condition $\mathbf{X}(0)=\mathbf{X}$ is:
$\partial_{t}\mathbf{X}(t)=-\mathbf{\mathcal{L}}_{\mathcal{B}}\mathbf{X}(t)\ ,\ \mathbf{X}(0):=\mathbf{X}\ .$
It is well known (see Hansen and Gebhart [16]) that the solution to this type of equation can be written using matrix exponentiation as:
$\mathbf{X}(t)=\exp(-t\mathbf{\mathcal{L}}_{\mathcal{B}})\mathbf{X}(0)$
We call the operator $\mathcal{H}_{\mathcal{B}}(t):=\exp(-t\mathbf{\mathcal{L}}_{\mathcal{B}})\in\mathbb{R}^{nd\times nd}$ the bundle heat kernel, which is defined as
$\mathcal{H}_{\mathcal{B}}(t):=\lim_{K\to\infty}\sum_{k=0}^{K}\frac{\left(-t\mathcal{L}_{\mathcal{B}}\right)^{k}}{k!}.$
Computing the heat kernel is necessary to solve the heat equation. An exact solution can be computed using spectral methods. For small $t$, one can instead consider the truncated Taylor expansion centered at 00, which amounts to fixing a $K$ in the above equation. Bodnar et al. [13] instead approximates the solution using the Euler discretization with unit time step of the heat equation.
However, all these methods pose a challenge when the bundle structure is learned as in Bodnar et al. [13], since the heat kernel has to be recomputed after every gradient update of the bundle structure. This high computational overhead limits the usability of Sheaf Neural Networks in applications.
Figure 2: Comparison of different Laplacians, from left to right: standard Laplacian, general bundle Laplacian, flat bundle Laplacian. Flat bundle Laplacian is amenable since it can be factored into a synchronization step, a standard Laplacian, and a desynchronization step.
@@
where $\mathbf{O}\in\mathbb{R}^{nd\times nd}$ is block diagonal with $v$-th block being $\mathbf{O}_{v}$. We call the matrices $\mathbf{O}$ and $\mathbf{O}^{T}$ synchronization and desynchronization, respectively. We compare different Laplacians in Figure 2. This factorization avoids the $\mathcal{O}\left(d^{3}|\mathsf{E}|\right)$ cost of computing the restriction map over each edge. Additionally, it allows to cast the bundle heat equation into a standard graph heat equation:
@@
This reduces the computation of the bundle heat kernel to that of the cheaper graph heat kernel, an operator that does not change depending on the bundle and can, therefore, also be pre-computed.
Our model. The BuNN layer occurs in four steps, as illustrated in Figure 1 in which steps 1, 2, 3, and 4 correspond to the Equations below of the same number. First, the bundle maps $\mathbf{O}_{v}$ are computed using a neural network $\boldsymbol{\phi}$ and the use of Householder reflections [17] or direct parameterization when $d=2$. Second, an encoder step updates the initial node signals via a matrix $\mathbf{W}\in\mathbb{R}^{d\times d}$, and bias $\mathbf{b}\in\mathbb{R}^{d}$. Next, the features are diffused over the learned vector bundle. Finally, a non-linearity $\sigma$ is applied. We summarize the steps in the following equations:
$\mathbf{O}^{(\ell)}_{v}$ $:=\boldsymbol{\phi}^{\left(\ell\right)}(\mathsf{G},\mathbf{X^{(\ell)}},v)\ \ \ \forall v\in\mathsf{V}$ (1)
@@
The diffusion time $t$ in Equation 3 is a hyperparameter determining the scale at which messages are diffused. For the case of small $t$, we approximate the heat kernel via its truncated Taylor series of degree $K$, and for large $t$, we use the exact solution using spectral methods. For simplicity of exposition, the steps above describe an update given a single bundle (i.e., $c=d$), meaning that $\mathbf{x}_{v}^{(\ell)}\in\mathbb{R}^{d}$. In general, we allow for multiple bundles and vector field channels, which we detail in Appendix E. We note that both Equations 2 and 3 are linear (or affine), and the non-linearities lie in 1 and 4. Equation 2 may be interpreted as a bundle-aware encoder, while Equation 3 is the message diffusion step guided by the heat kernel.
@@
We note that without the bundle structure, Equation 3 would converge exponentially fast to constant node representations over the graph (e.g., Theorem $1$ in [4]), potentially leading to over-smoothing. Accordingly, the bundle is crucial in this formulation to prevent node features from converging to constant. We note that this is a limitation of existing diffusion-based GNNs [18, 19].
@@
Link to Graph Convolutional Networks. It is possible to derive Graph Convolutional Networks (GCNs) [15] as an approximation of BuNNs operating over a trivial bundle. Setting $t=1$ Equation 3 becomes $\mathbf{Z}^{(l+1)}=\exp(-\mathcal{L}_{\mathsf{G}})\mathbf{H}^{(l)}$. The approximation $\exp(-\mathcal{L}_{\mathsf{G}})\approx 1-\mathcal{L}_{\mathsf{G}}$ gives the update $\mathbf{Z}^{(l+1)}=(1-\mathcal{L}_{\mathsf{G}})\mathbf{H}^{(l)}=\mathbf{A}_{\mathsf{G}}\mathbf{H}^{(l)}$ recovering the GCN update. Section 4.1 shows that BuNNs can avoid over-smoothing, highlighting an important difference between GCNs and BuNNs.
@@
Comparison with Sheaf Neural Networks. Flat vector bundles are a special case of cellular sheaves [14, 13], meaning that our model has close connections to Sheaf Neural Networks (SNNs) [16, 13, 20, 21, 22], with the closest relative being the Neural Sheaf Diffusion (NSD) model from Bodnar et al. [13]. BuNNs distinguish themselves from NSD in several ways. First, NSD approximates the heat equation using a time-discretized solution to the heat equation, which results in a standard message passing algorithm. In contrast, the direct use of the heat kernel allows BuNNs to break away from the explicit message-passing paradigm. Secondly, the use of flat bundles increase scalability since the bundle maps are computed at the node level. Additionally, flat bundles guarantees path independence, a requirement for the theory in Bodnar et al. [13] to hold, often not satisfied for general sheaf constructions such as ones used in NSD. Thirdly, we allow $\phi$ to be any GNN while NSD restricts $\phi$ to be an MLP. We found that incorporating the graph structure in $\phi$ improved the experimental results. The update in Equation 2 is also different to NSD, and is necessary to prove our main theoretical result, Theorem 5.3. We provide experimental comparisons to NSDs in the experimental section and show that BuNNs significantly outperform their sheaf counterparts. We provide a summarized comparison between GCNs, SNNs, and BuNNs in Table 1.
@@
Table 1: Comparison between GCN, SNN, and BuNN in terms of propagation and message type, as well as their capability of mitigating issues related to GNNs.
@@
No under-reaching
No over-squashing
No over-smoothing

### v2

@@ Abstract
The dominant paradigm for learning on graphs is message passing. Despite being a strong inductive bias, the local message passing mechanism faces challenges such as over-smoothing, over-squashing, and limited expressivity. To address these issues, we introduce Bundle Neural Networks (BuNNs), a novel graph neural network architecture that operates via message diffusion on flat vector bundles — geometrically inspired structures that assign to each node a vector space and an orthogonal map. A BuNN layer evolves node features through a diffusion-type partial differential equation, where its discrete form acts as a special case of the recently introduced Sheaf Neural Network (SNN), effectively alleviating over-smoothing. The continuous nature of message diffusion enables BuNNs to operate at larger scales, reducing over-squashing. We establish the universality of BuNNs in approximating feature transformations on infinite families of graphs with injective positional encodings, marking the first positive expressivity result of its kind. We support our claims with formal analysis and synthetic experiments. Empirically, BuNNs perform strongly on heterophilic and long-range tasks, which demonstrates their robustness on a diverse range of challenging real-world tasks.
@@ 1 Introduction
Graph Neural Networks (GNNs) (Sperduti, 1993; Scarselli et al., 2009; Defferrard et al., 2016) are widely adopted machine learning models designed to operate over graph structures, with successes in diverse applications such as drug discovery (Stokes et al., 2020), traffic forecasting (Derrow-Pinion et al., 2021), and recommender systems (Fan et al., 2019). Most GNNs are Message Passing Neural Networks (MPNNs) (Gilmer et al., 2017), where nodes exchange messages with immediate neighbours. While effective, MPNNs face critical challenges such as over-smoothing (Li et al., 2018; Oono & Suzuki, 2020; Cai & Wang, 2020), over-squashing (Alon & Yahav, 2021; Topping et al., 2022; Di Giovanni et al., 2023), and limited expressivity (Xu et al., 2019b; Morris et al., 2019).
Figure 1: Local message passing on graphs versus global message diffusion on bundles.
Over-smoothing occurs when node features become indistinguishable as the depth of the MPNN increases, a problem linked to the stable states of the heat equation on graphs (Cai & Wang, 2020). While Sheaf Neural Networks (Bodnar et al., 2022) address this by enriching the graph with a sheaf structure that assigns linear maps to edges and results in richer stable states of the corresponding heat equation, they remain MPNNs and inherit other limitations such as over-squashing, which restricts the amount of information that can be transmitted between distant nodes.
We propose Bundle Neural Networks (BuNNs), a new type of global GNN that operates over flat vector bundles – structures analogous to connections on flat Riemannian manifolds that augment the graph by assigning to each node a vector space and an orthogonal map. BuNNs do not perform ‘explicit’ message passing through multiple steps of information exchange between neighboring nodes, but instead operate via message diffusion. Each layer involves a node update step, and a diffusion step evolving the features according to a vector diffusion PDE as in Singer & Wu (2012). The resulting architecture enjoys the desirable properties of Sheaf Neural Networks, in that they can avoid over-smoothing, but are global models that can operate at larger scales of the graph to tackle over-squashing, and it achieves better performance on a range of benchmark datasets. Additionally, we prove that equipped with injective positional encodings, BuNNs are compact uniform approximators, a new type of universality result for feature transformation approximation.
@@
  • We prove that the diffusion process can mitigate over-smoothing and over-squashing (Section 4), and support these claims with novel synthetic experiments (Section 6.1).
  • We prove that, with injective positional encodings, BuNNs are compact uniform universal approximators. To the best of our knowledge, this is the first of such results (Section 5).
  • We show that BuNNs perform well on heterophilic and long-range tasks, for instance, achieving a new state-of-the-art result on the Peptides-func dataset (Section 6.2).
@@ 2 Background
Graphs. Let $\mathsf{G}=(\mathsf{V},\mathsf{E})$ be an undirected graph on $n=|\mathsf{V}|$ nodes with edges $\mathsf{E}$. We represent the edges via an adjacency matrix $\mathbf{A}\in\mathbb{R}^{n\times n}$ where the entry $\mathbf{A}_{uv}$ for $u,v\in\mathsf{V}$ is $1$ if the edge $(u,v)\in\mathsf{E}$ and 00 otherwise. Let $\mathbf{D}$ be the diagonal degree matrix with entry $\mathbf{D}_{vv}=d_{v}$ equal to the degree of $v$. The graph Laplacian is defined as $\mathbf{L}:=\mathbf{D}-\mathbf{A}$ and the random walk normalized graph Laplacian is defined as $\mathbf{\mathcal{L}}:=\mathbf{I}-\mathbf{D^{-1}A}$. We assume that at each node $v\in\mathsf{V}$ we are given a $c$-dimensional signal (or node feature) $\mathbf{x}_{v}\in\mathbb{R}^{c}$ and group such signals into a matrix $\mathbf{X}\in\mathbb{R}^{n\times c}$.
@@
GNNs and feature transformations. A feature transformation on a graph $\mathsf{G}$ is a permutation-equivariant map $f_{\mathsf{G}}:\mathbb{R}^{n\times c_{1}}\rightarrow\mathbb{R}^{n\times c_{2}}$ that transforms the node signals. A $\operatorname{GNN}_{\bm{\Theta}}$ is a (continuous) map parameterized by $\bm{\Theta}$ that takes as input a graph alongside node signals $\mathsf{G}=(\mathsf{V},\mathsf{E},\mathbf{X})$ and outputs a transformed signal $(\mathsf{V},\mathsf{E},\mathbf{X^{\prime}})$. A GNN on a graph $\mathsf{G}$ is therefore a feature transformation $\operatorname{GNN}_{\Theta}:\mathbb{R}^{n\times c_{1}}\rightarrow\mathbb{R}^{n\times c_{2}}$. Given a collection of graphs $\mathcal{G}$, a feature transformation $F$ on $\mathcal{G}$ is an assignment of every graph $\mathsf{G}\in\mathcal{G}$ to a feature transformation $F_{G}:\mathbb{R}^{n_{G}\times c_{1}}\rightarrow\mathbb{R}^{n_{G}\times c_{2}}$. The set of continuous feature transformations over a collection of graphs in $\mathcal{G}$ is denoted $\mathcal{C}\left(\mathcal{G},\mathbb{R}^{c_{1}},\mathbb{R}^{c_{2}}\right)$.
@@
Cellular sheaves. A cellular sheaf (Curry, 2014) $\left(\mathcal{F},\mathsf{G}\right)$ over an undirected graph $\mathsf{G}=\left(\mathsf{V},\mathsf{E}\right)$ augments $\mathsf{G}$ by attaching to each node $v$ and edge $e$ a vector space space called stalks and denoted by $\mathcal{F}(v)$ and $\mathcal{F}(e)$, usually the stalks are copies of $\mathbb{R}^{d}$ for some $d$. Additionally, every incident node-edge pair $v\trianglelefteq e$ gets assigned a linear map between stalks called restriction maps and denoted $\mathcal{F}_{v\trianglelefteq e}:\mathcal{F}(v)\rightarrow\mathcal{F}(e)$. Given two nodes $v$ and $u$ connected by an edge $(v,\ u)$, we can transport a vector $\mathbf{x}_{v}\in\mathcal{F}(v)$ from $v$ to $u$ by first mapping it to the stalk at $e=(v,u)$ using $\mathcal{F}_{v\trianglelefteq e}$, and mapping it to $\mathcal{F}(u)$ using the transpose $\mathcal{F}_{u\trianglelefteq e}^{T}$. As a generalization of the graph adjacency matrix, the sheaf adjacency matrix $\mathbf{A}_{\mathcal{F}}\in\mathbb{R}^{nd\times nd}$ is defined as a block matrix in which each $d\times d$ block $(\mathbf{A}_{\mathcal{F}})_{uv}$ is $\mathcal{F}_{u\trianglelefteq e}^{T}\mathcal{F}_{v\trianglelefteq e}$ if there is an edge between $u$ and $v$ and $\mathbf{0}_{d\times d}$ otherwise. Similary, we define the block diagonal degree matrix $\mathbf{D}_{\mathcal{F}}\in\mathbb{R}^{nd\times nd}$ as $\left(\mathbf{D}_{\mathcal{F}}\right)_{vv}:=d_{v}\mathbf{I}_{d\times d}$, and the sheaf Laplacian is $\mathbf{L}_{\mathcal{F}}:=\mathbf{D_{\mathcal{F}}}-\mathbf{A}_{\mathcal{F}}$. These matrices act as bundle generalizations of their well-known standard graph counterparts and we recover such matrices when $\mathcal{F}(v)\cong\mathbb{R}$ and $\mathcal{F}_{v\trianglelefteq e}{v}=1$ for all $v\in\mathsf{V}$ and $e\in\mathsf{E}$.
@@
Vector bundles. When restriction maps are orthogonal, we call the sheaf a vector bundle, a structure analogous to connections on Riemannian manifolds. For this reason, the sheaf Laplacian also takes the name connection Laplacian (Singer & Wu, 2012). The product $\mathcal{F}_{u\trianglelefteq e}^{T}\mathcal{F}_{v\trianglelefteq e}$ is then also orthogonal and is denoted $\mathbf{O}_{uv}$ referring to the transformation a vector undergoes when moved across a manifold via parallel transport. In this case we denote the node-stalk at $v$ by $\mathcal{B}(v)$, the bundle-adjacency by $\mathbf{A}_{\mathcal{B}}$ and the bundle Laplacian $\mathbf{L}_{\mathcal{B}}$, and its normalized version $\mathbf{\mathcal{L}}_{\mathcal{B}}:=\mathbf{I}_{dn\times dn}-\mathbf{D}_{\mathcal{B}}^{-1}\mathbf{A}_{\mathcal{B}}$.
@@ 3 Bundle Neural Networks
In this section, we derive BuNN from heat diffusion equations over flat vector bundles. We then discuss its relationship to GCN (Kipf & Welling, 2017) and NSD (Bodnar et al., 2022). We provide algorithmic implementations and additional practical details in the Appendix (Section E).
@@
The gradient of the Dirichlet energy $\nabla_{\mathbf{X}}\mathcal{E}_{\mathcal{B}}(\mathbf{X})$ is the random-walk Laplacian $\mathbf{\mathcal{L}}_{\mathcal{B}}$. We can write down a heat diffusion equation over a vector bundle as a gradient flow, whose evolution equation with initial condition $\mathbf{X}(0)=\mathbf{X}$ satisfies $\partial_{t}\mathbf{X}(t)=-\mathbf{\mathcal{L}}_{\mathcal{B}}\mathbf{X}(t)$. The solution to this equation can be written using matrix exponentiation as $\mathbf{X}(t)=\exp(-t\mathbf{\mathcal{L}}_{\mathcal{B}})\mathbf{X}(0)$ (e.g. Hansen & Gebhart (2020)). We call the operator $\mathcal{H}_{\mathcal{B}}(t):=\exp(-t\mathbf{\mathcal{L}}_{\mathcal{B}})\in\mathbb{R}^{nd\times nd}$ the bundle heat kernel, which is defined as:
$\mathcal{H}_{\mathcal{B}}(t)=\lim_{K\to\infty}\sum_{k=0}^{K}\frac{\left(-t\mathcal{L}_{\mathcal{B}}\right)^{k}}{k!}.$
Computing the heat kernel is necessary to solve the heat equation. An exact solution can be computed using spectral methods. For small $t$, one can instead consider the truncated Taylor expansion centered at 00, which amounts to fixing a $K$ in the above equation. Bodnar et al. (2022) instead approximates the solution using the Euler discretization of the heat equation with unit time step.
However, all these methods pose a challenge when the bundle structure is learned as in Bodnar et al. (2022), since the heat kernel has to be recomputed after every gradient update of the bundle structure. This high computational overhead limits the usability of Sheaf Neural Networks in applications.
Figure 2: Comparison of different Laplacian and their actions on signals.
@@
where $\mathbf{O}\in\mathbb{R}^{nd\times nd}$ is block diagonal with $v$-th block being $\mathbf{O}_{v}$. We call the matrices $\mathbf{O}$ and $\mathbf{O}^{T}$ synchronization and desynchronization, respectively. We compare different Laplacians in Figure 2. This factorization avoids the $\mathcal{O}\left(d^{3}|\mathsf{E}|\right)$ cost of computing the restriction map over each edge. Additionally, Lemma 3.1 shows that it allows to cast the bundle heat equation into a standard graph heat equation. This reduces the computation of the bundle heat kernel to that of the cheaper graph heat kernel, an operator that does not change depending on the bundle and can, therefore, also be pre-computed.
@@
The model. The BuNN layer occurs in four steps, as illustrated in Figure 3. First, the bundle maps $\mathbf{O}_{v}$ are computed using a neural network $\bm{\phi}$, the graph $\mathsf{G}$, positional encodings $\mathbf{P}\in\mathbb{R}^{n\times f}$ and the use of Householder reflections (Householder, 1958) or direct parameterization of the orthogonal group when $d=2$. Second, an encoder step updates the node signals via a learnable matrix $\mathbf{W}\in\mathbb{R}^{d\times d}$, and bias $\mathbf{b}\in\mathbb{R}^{d}$. Next, the features are diffused over the learned vector bundle using the heat kernel. Finally, a non-linearity $\sigma$ is applied. We summarize the steps in the following equations:
$\mathbf{O}^{(\ell)}_{v}$ $:=\bm{\phi}^{\left(\ell\right)}(\mathsf{G},\mathbf{P},\mathbf{X^{(\ell)}},v)\ \ \ \forall v\in\mathsf{V}$ (1)
@@
The diffusion time $t$ in Equation 3 is a hyperparameter determining the scale at which messages are diffused. For the case of small $t$, we approximate the heat kernel via its truncated Taylor series of degree $K$, and for large $t$, we use spectral methods. For simplicity of exposition, the steps above describe an update given a single bundle (i.e., $c=d$), meaning that $\mathbf{x}_{v}\in\mathbb{R}^{d}$. In general, we allow multiple bundles and vector field channels (Appendix E). Note that Equations 1, 2, and 3 are linear (or affine), and the non-linearities lie in 4. Equation 2 may be interpreted as a bundle-aware encoder, while Equation 3 is the message diffusion step guided by the heat kernel.
@@
Without the bundle structure, Equation 3 would converge exponentially fast to constant node representations over the graph (e.g. Theorem $1$ in Li et al. (2018)), potentially leading to over-smoothing. This is a limitation of existing diffusion-based GNNs (Xu et al., 2019a; Zhao et al., 2021). Accordingly, the bundle is crucial in this formulation to prevent node features from collapsing.
@@
Link to Graph Convolutional Networks. It is possible to derive Graph Convolutional Networks (GCNs) (Kipf & Welling, 2017) as an approximation of BuNNs operating over a trivial bundle. Setting $t=1$ Equation 3 becomes $\mathbf{Z}^{(l+1)}=\exp(-\mathcal{L}_{\mathsf{G}})\mathbf{H}^{(l)}$. The approximation $\exp(-\mathcal{L}_{\mathsf{G}})\approx\mathbf{I}-\mathcal{L}_{\mathsf{G}}$ gives the update $\mathbf{Z}^{(l+1)}=(1-\mathcal{L}_{\mathsf{G}})\mathbf{H}^{(l)}=\mathbf{A}_{\mathsf{G}}\mathbf{H}^{(l)}$ recovering the GCN update.
@@
Comparison with Sheaf Neural Networks. Flat vector bundles are a special case of cellular sheaves (Curry, 2014; Bodnar et al., 2022), meaning that our model has close connections to Sheaf Neural Networks (SNNs) (Hansen & Gebhart, 2020; Bodnar et al., 2022; Barbero et al., 2022b; a). While most SNNs operate on fixed sheaves (Hansen & Gebhart, 2020; Barbero et al., 2022a; Battiloro et al., 2023), we focus on learning sheaves as in Neural Sheaf Diffusion (NSD) from Bodnar et al. (2022). BuNNs distinguish themselves from NSD in several ways. First, NSD approximates the heat equation using a time-discretized solution to the heat equation, which results in a standard message passing algorithm. In contrast, the direct use of the heat kernel allows BuNNs to break away from the explicit message-passing paradigm. This is in line with Battiloro et al. (2024) who define convolutions on manifolds using the heat kernel of fixed sheaves. Secondly, the use of flat bundles increases scalability since the bundle maps are computed at the node level. Additionally, flat bundles guarantees path independence, a requirement for the theory on the long time limit of NSD in Bodnar et al. (2022) to hold, often not satisfied for general sheaf constructions such as ones used in NSD. Thirdly, we allow $\phi$ to be any GNN while NSD restricts $\phi$ to be an MLP. We found that incorporating the graph structure in $\phi$ improved the experimental results. The update in Equation 2 is also different to NSD in how the $\mathbf{W}$ and $\mathbf{b}$ are applied, and is necessary to prove our main theoretical result, Theorem 5.3. Additionally, Bodnar et al. (2022) experiments with general restriction maps as opposed to restricting them to orthogonal maps, and find that in practice orthogonal restrictions perform better, hence we consider orthogonal restriction maps. We provide experimental comparisons to NSDs in the experimental section and show that BuNNs significantly outperform their sheaf counterparts. We provide a summarized comparison between GCNs, SNNs, and BuNNs in Table 1.
Comparison with other methods. Another paradigm for learning on graphs is using graph transformers (GT), where every nodes communicate to each other through the use of self-attention. When the graph transformer is fully connected, all nodes communicate and therefore GT should not suffer from under-reaching or over-squashing, but might suffer from over-smoothing (Dovonon et al., 2024). When not fully connected, they might suffer from over-squashing (Barbero et al., 2024). Other paradigms such as Implicit Graph Neural Networks Fu et al. (2023) are also designed to tackle under-reaching and long-range dependencies while empirically not suffering from over-smoothing.
@@
Table 1: Comparison between models in terms of message type and capability of mitigating issues related to GNNs.
@@
No
under-reaching
Alleviates
over-squashing
Alleviates
over-smoothing